OpenStudy (anonymous):

5+square root of 2x+1=10

OpenStudy (amoodarya):

\[5+\sqrt{2x+1}=10\] is it ?

OpenStudy (amoodarya):

\[\sqrt{2x+1}=10 -5 \rightarrow \sqrt{2x+1}=5 \rightarrow (\sqrt{2x+1}=5)^2\]

OpenStudy (amoodarya):

then simplify and find x then check x in original equation